Please show work/explain:

1)

The molar mass of succinic acid is 118.09g/mol. Suppose a 1.1519g sample of succinic acid is titrated with 25.95mL of 0.9913M NaOH solution. Is Succinic acid monoprotic/diprotic/triprotic or tetraprotic and why, show work.

2)  

2C2H6(g)+7o2(g)→4co2(g)+6h2o(g) Standard Enthalpy is -3120.6kj/mol

What mass of ethane must be combusted in order to heat 366g of water from 25.0C to 100.0C. ASsume the heat transfer is 100% efficient?

3)

2HCl + caco3 --> H20 + CO2 + CaCl2. Assuming the rxn takes place at a constant temperature and pressure which of the following statements is true and why?

a)the system does work on the surroundings

b) the surroundings does work on the system
Answer #1

1) The molar mass of succinic acid given in question is 118.09g/mol.

1.1519g of succinic acid is titrated with 25.95mL of 0.9913M NaOH solution.

The formula of succinic acid is C4H6O4 and the structure is:

HOOC-CH2-CH2-COOH.

The structure have two acidic hydrogens or we can say that the succinic acid has two replacable hydrogens. Therefore , Succinic acid will be diprotic.

2) Standard Enthalpy is -3120.6 kj/mol

mass of ethane = ?

mass of water = 366g

temperature change from 25oC to 100oC

Efficiency of heat transfer = 100 %

The reaction involved is as follows:

2C2H6 + 7O2 = 4CO2 + 6H2O

For producing 6 molecules of water  = 366 g of water is needed

Then, for 1 molecule = 366/6

Hence, for combustion of 2 moles of ethane, the mass of ethane is needed = (366 * 2) / 6

= 122 g.

The mass of ethane is needed for combustion is equal to 122g.

3)  

CaCO3 + 2HCl = CaCl2 + CO2 + H2O

  • The reaction between hydrochloric acid and calcium carbonate (which is also known as marble chips) is endothermic in nature because the change in enthalpy is positive. that is as follows:
  • Change in Enthalpy: ΔH = +79.9kJ.
  • In this reaction the physical change is endothermic which means the heat is giving by the surrounding to the system because of this change the temperature of the surrounding become decreases.
  • Hence, The second statement is true that is the surrounding does work on the system.